# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

### v1

@@ C R C R
Official Baseline (Park et al., 2018) 7.0 12.0 6.9 11.9
XE (Wang et al., 2018) 7.2 11.5 - -
AREL (Wang et al., 2018) 7.3 11.4 - -
@@
(a) Results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE.
@@
Official Baseline (Park et al., 2018) 4.015
@@
(b) Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better.
@@
(c) Human evaluation results on VIST. Higher is better.
Figure 9. TAPM evaluation results. Adapted from (Yu et al., 2021).
6.4. Combining LLMs (GPT-2) and Character Features
@@
Hong et al. (Hong et al., 2023) proposed CharGrid (Character-Grid Transformer), a transformer model that integrates diverse input features, from global image attributes to character-specific details. It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s (Liu et al., 2021a), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing. The trainable image and character encoders process the image feature inputs. The character grid is flattened and fed to a FC layer. The transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained GPT-2 tokenizer. The model architecture is shown in Fig. 10.
@@
Figure 10. CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from (Hong et al., 2023).
@@
During training, model parameters, excluding the vision backbone, are randomly initialized. The training objective involves maximizing the likelihood of image sequence-story pairs through backpropagation. Nucleus sampling (Holtzman et al., 2019) is employed for validation, and the METEOR score is used for evaluation.
@@
Table 5. Results using different input features on the test set of VWP. Global features are always included. “+ obj” and “+ char” represent that object and character features are also included, respectively. BLEU (B), METEOR (M), ROUGE-L (R-L), and CIDEr (C) values are averages of three runs with different random seeds. Adapted from (Hong et al., 2023).
Model Grammaticality Coherence Visual Groundedness Diversity
@@
Table 6. Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality, Coherence, Visually Groundedness, and Diversity. The numbers are percentages. * means p-value $<$ 0.05. ** means p-value $<$ 0.01. Adapted from (Hong et al., 2023).
6.5. Visual Story Generation using Graphs for Event Ratiocination
Zheng et al. (Zheng et al., 2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters. The work addresses several downstream tasks, including VSG. The work introduces a framework to improve the integration and interpretation of visual and textual data, using hypergraphs to address challenges in multimodal interactions and temporal dynamics. Figure 11 shows this architecture.
Figure 11. Architecture of HEGR by Zheng et al. Adapted from (Zheng et al., 2021).
The method begins with feature extraction. Visual features are extracted using Mask R-CNN (He et al., 2017) to detect objects and their spatial characteristics within images. Textual features are extracted from associated descriptions or captions using BERT (Devlin et al., 2019), capturing semantic information. In the second phase, semantic graphs are constructed for each modality: nodes represent detected features or semantic tokens, while edges depict relationships within the same modality. To capture dependencies between modalities and across various temporal segments, hypergraphs are employed, linking semantic graphs based on contextual and temporal relevance, forming a high-dimensional interaction map. In the third phase, graph convolution is used to update node representations within each graph, allowing contextual information aggregation through learned adjacency matrices. Hyper-graph convolution extends this to include higher-order relationships across multiple graphs and temporal instances. In the fourth phase, a self-attention mechanism refines graph representations by dynamically adjusting the significance of nodes within and across modalities. This process focuses on the most relevant features for narrative generation. Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART (Lewis et al., 2019).
@@ 6.6. Discussion
Specifically, the first three techniques, presented in Sections 6.16.2, and 6.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation. The methods detailed in Sections 6.1 and 6.3, offer comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively. Notably, the latter technique surpasses the former, which can be attributed to its transitional training approach.
@@
Table 7 summarizes the performance of the methods described above, along with other methods from the literature that contributed to the field of VSG. The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN (Ren et al., 2015) and Mask-RCNN (He et al., 2017), and LLMs like GPT-2 (Radford et al., 2019) and BART (Lewis et al., 2019). The first two methods, Park et al. (2015) (Park and Kim, 2015) and Park et al. (2019) (Park et al., 2018) are not comparable to the other methods as they do not use the VIST dataset. The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) (Hong et al., 2023) that is evaluated on the VWP dataset. The authors of CharGrid trained and evaluated the performance of TAPM on the VWP (Hong et al., 2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset. It is however not possible to know if CharGrid would perform better than HEGR on the VIST dataset without further experiments.
@@
Park et al. (2015) (Park and Kim, 2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
Park et al. (2019) (Park et al., 2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
Huang et al. (2015) (Huang et al., 2016) CNN, GRU GRU VIST - - 31.4
Xu et al. (2015) (Xu et al., 2015) VGGNet LSTM VIST - 28.94 32.98
Yu et al. (2017) (Yu et al., 2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
CST (2018) (Gonzalez-Rico and Pineda, 2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
AREL (2018) (Wang et al., 2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
HSRL (2019) (Huang et al., 2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
GLACNet (2019) (Kim et al., 2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
StoryAnchor (2020) (Modi and Parde, 2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
INET (2020) (Jung et al., 2020) CNN, GRU GRU VIST 14.7 29.7 35.5
AOG-LSTM (2023) (Liu et al., 2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
CoVS (2023) (Gu et al., 2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
TAPM (2021) (Yu et al., 2021) 1 Faster R-CNN, ResNet-152 GPT-2 VIST - 33.1 37.2
HEGR (2021) (Zheng et al., 2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
TAPM (2021) (Hong et al., 2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
CharGrid (2023) (Hong et al., 2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
@@
Table 7. Analysis of Visual Story Generation works. The columns “B-4”, “R-L”, and “M” represent the BLEU-4, ROUGE-L, and METEOR scores, respectively. All works, except Huang et al. (2015) (Huang et al., 2016) and CST (2018) (Gonzalez-Rico and Pineda, 2018), use attention mechanisms. All works use images as input, except ActivityNet (Fabian Caba Heilbron and Niebles, 2015) (it uses videos). The mark 1 indicates works that used extra training data (direct comparison may not be fair). The mark 2 indicated works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
7. Conclusions

### v2

@@ C R C R
Baseline Park et al. (2018) 7.0 12.0 6.9 11.9
XE Wang et al. (2018) 7.2 11.5 - -
AREL Wang et al. (2018) 7.3 11.4 - -
@@
Table 6: Quantitative results on LSMDC 2019 public and blind test set. C stands for CIDEr and R for ROUGE. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021).
@@
Official Baseline Park et al. (2018) 4.015
@@
Table 7: Human evaluation results on the LSMDC 2019 blind test set according to a Likert scale from 5 (worst) to 1 (best), where lower is better. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021).
@@
Table 8: Human evaluation results on VIST. Higher is better. Green indicates the highest score, red the lowest. Adapted from Yu et al. (2021)
5.4 Interactive and Creative Visual Storytelling
Building upon the foundational CNN and RNN approaches, interactive and creative visual storytelling represents an important evolution in the field, introducing intermediate approaches that bridge visual comprehension with narrative creativity. These methods combine elements of image captioning and VQA while introducing narrative structure and creative interpretation, addressing limitations of purely technical approaches.
Lukin et al. Lukin et al. (2018) introduced a three-module pipeline for creative visual storytelling that systematically approaches the challenge of generating narratives from visual content. Their pipeline consists of object identification, single-image inferencing, and multi-image narration, that serve as a preliminary design for building a creative visual storyteller. The approach defines computational creative visual storytelling as one with the ability to alter the telling of a story along three aspects: to speak about different environments, to produce variations based on narrative goals, and to adapt the narrative to the audience. This modular approach demonstrates how the foundational capabilities of image captioning and VQA can be systematically extended and combined to support more complex narrative generation tasks. The pipeline explicitly separates visual understanding from narrative construction, allowing for more systematic development and evaluation of each component while maintaining the overall coherence required for storytelling.
Interactive visual storytelling has also explored user engagement and personalization in narrative generation. Halperin and Lukin Halperin and Lukin (2023) examined how creative visual storytelling can serve as an anthology for narrative intelligence, investigating the intersection of human creativity and automated story generation through an analysis of 100 visual stories from authors who participated in a systematic creative process of improvised story building based on image sequences. Their work on surreal visual storytelling Halperin and Lukin (2024) further explores how visual narrative systems can handle ambiguous or dreamlike imagery, investigating AI “hallucination” by stress-testing a visual storytelling algorithm with different visual and textual inputs designed to probe dream logic inspired by cinematic surrealism. These approaches highlight the importance of modularity in visual storytelling systems, where different components can be optimized independently while contributing to the overall narrative coherence. The integration of creative elements with systematic visual analysis provides a foundation for more sophisticated VSG systems that can balance factual visual description with imaginative narrative construction.
5.5 Knowledge-Enhanced Visual Storytelling
A challenge in VSG is the tendency of end-to-end approaches to produce monotonous stories with repetitive text and limited lexical diversity. This limitation arises because existing approaches are constrained by the vocabulary and knowledge available in single training datasets. To address this challenge, researchers have explored knowledge enhanced approaches that leverage external resources to enrich the story generation process.
Hsu et al. Hsu et al. (2020) introduced KG-Story, a three-stage framework that allows VSG systems to take advantage of external knowledge graphs to produce more diverse stories. The framework implements a distill-enrich-generate approach: first distilling a set of representative words from input prompts, then enriching the word set using external knowledge graphs, and finally generating stories based on the enriched word set. This framework allows the use of external resources not only for the enrichment phase, but also for the distillation and generation phases. The KG-Story framework operates through three distinct stages. In stage 1, an image-to-term model distills representative terms from each input image, creating conceptual representations that capture the essential elements of the visual content. In stage 2, external knowledge graphs are utilized to identify possible connections between the extracted term sets from different images, generating enriched term paths that capture relationships and associations not explicitly visible in the images. In stage 3, a Transformer architecture transforms these term paths into coherent stories, incorporating techniques such as length difference positional encoding and repetition penalties to improve narrative quality. Evaluation results demonstrate that stories generated by KG-Story are on average ranked better than previous state-of-the-art systems in human ranking evaluations. The approach successfully addresses the vocabulary limitations of traditional end-to-end methods while maintaining narrative coherence.
Interactive approaches to knowledge enhanced storytelling have also been explored. Hsu et al. Hsu et al. (2019) introduced Dixit, an interactive visual storytelling system that allows users to iteratively compose stories through term manipulation. The system extracts text terms describing objects and actions from photos, then allows users to add new terms or remove existing ones before generating stories based on these modified term sets. Behind the scenes, Dixit uses an LSTM model trained on image caption data to distill terms from images and utilizes a Transformer decoder to compose context coherent stories. This approach opens up possibilities for interpretable and controllable visual storytelling, allowing users to understand the story formation rationale and to intervene in the generation process.
These knowledge enhanced approaches represent an important advancement in VSG, demonstrating how external knowledge resources can be systematically integrated to overcome the limitations of purely data-driven methods. By explicitly modeling the relationship between visual content and broader conceptual knowledge, these systems can generate more diverse, engaging, and contextually rich narratives while maintaining controllability and interpretability.
5.6 Iterative and Planning-Based Approaches
Traditional VSG models typically generate stories in a single forward pass. However, creative writers use their knowledge and worldview to put disjointed elements together to form a coherent storyline, and work and rework iteratively toward perfection. VSG models, however, make poor use of external knowledge and iterative generation when attempting to create stories. This observation has motivated the development of iterative and planning-based approaches that more closely mirror human creative writing processes. Hsu et al. Hsu et al. (2021) introduced PR-VIST (Plot and Rework Visual Storytelling), a framework that represents the input image sequence as a story graph in which it finds the best path to form a storyline. PR-VIST then takes this path and learns to generate the final story via a re-evaluating training process. The framework draws inspiration from the human creative writing process, which involves plotting (planning the overall narrative structure) and reworking (iteratively refining the story content).
The PR-VIST framework operates through two main stages that correspond to the plotting and reworking phases of creative writing. In stage 1 (story plotting), the system constructs a story graph from the input image sequence using external knowledge graphs, including VIST and Visual Genome knowledge graphs. A storyline predictor model identifies the best path through this graph to form a coherent storyline, creating a structured narrative plan that connects the images through meaningful conceptual relationships. This plotting stage essentially creates a roadmap for the story that will be generated. In stage 2 (story reworking), the framework generates the actual story text based on the predicted storyline path. The approach uses a length controlled Transformer that is first pre-trained on the ROC Story dataset and then fine-tuned on the VIST dataset with a discriminator component to encourage higher quality story generation. The re-evaluating training process allows the model to iteratively improve the story generation by learning from feedback about narrative quality and coherence.
Evaluation results demonstrate that this framework produces stories that are superior in terms of diversity, coherence, and humanness, per both automatic and human evaluations. An ablation study shows that both plotting and reworking contribute to the model’s superiority. The explicit separation of planning and generation phases allows for better control over narrative structure while maintaining the flexibility needed for creative story generation. This planning based approach represents an important advancement toward more human-like story generation, demonstrating how computational systems can benefit from explicitly modeling the structured creative processes that human writers naturally employ.
5.7 Combining LLMs (GPT-2) and Character Features
@@
Hong et al. Hong et al. (2023) proposed CharGrid (Character-Grid Transformer), a Transformer model that integrates diverse input features, from global image attributes to character-specific details. It takes a sequence of input tokens, including global image features obtained using Swin Transformer’s Liu et al. (2021), character features extracted by cropping character instances, a character grid for coherence assessment, and text features introduced incrementally during processing. The trainable image and character encoders process the image feature inputs. The character grid is flattened and fed to a FC layer. The Transformer module then processes these inputs, generating an output as a probability distribution over potential tokens using a pre-trained Generative Pre-trained Transformer (GPT)-2 tokenizer. The model architecture is shown in Fig. 3.
@@
Figure 3: CharGrid architecture: the blue boxes are pre-trained components where the parameters are fixed. Adapted from Hong et al. (2023).
@@
During training, model parameters, excluding the vision backbone, are randomly initialized. The training objective involves maximizing the likelihood of image sequence-story pairs through backpropagation. Nucleus sampling Holtzman et al. (2019) is employed for validation, and the METEOR score is used for evaluation.
@@
Table 9: Results using different input features on the test set of VWP. Global features are always included. “+ obj” and “+ char” represent that object and character features are also included, respectively. BLEU (B), METEOR (M), ROUGE-L (R-L), and CIDEr (C) values are averages of three runs with different random seeds. Green indicates the highest score, red the lowest. Adapted from Hong et al. (2023).
Model G C VG D
@@
Table 10: Human binary judgments in percentage of generated stories between TAPM and TAPM with character features (TAPM + char), TAPM + char and CharGrid on the test set of VWP across four criteria: Grammaticality (G), Coherence (C), Visual Groundedness (VG), and Diversity (D). The numbers are percentages. * means p-value << 0.05. ** means p-value << 0.01. Adapted from Hong et al. (2023).
5.8 Visual Story Generation using Graphs for Event Ratiocination
Zheng et al. Zheng et al. (2021) introduced Hypergraph-Enhanced Graph Reasoning (HEGR) for “visual event ratiocination”, a task that involves generating interpretations and narratives for events that precede, coincide with, or follow a visual scene, as well as discerning the intents of depicted characters. The work addresses several downstream tasks, including VSG. The work introduces a framework to improve the integration and interpretation of visual and textual data, using hypergraphs to address challenges in multimodal interactions and temporal dynamics. Fig. 4 shows this architecture.
Figure 4: Architecture of HEGR by Zheng et al. Adapted from Zheng et al. (2021).
The method begins with feature extraction: visual features are extracted using Mask R-CNN He et al. (2017) to detect objects and their spatial characteristics within images. Textual features are extracted from associated descriptions or captions using BERT Devlin et al. (2019), capturing semantic information. In the second phase, semantic graphs are constructed for each modality: nodes represent detected features or semantic tokens, while edges depict relationships within the same modality. To capture dependencies between modalities and across various temporal segments, hypergraphs are employed, linking semantic graphs based on contextual and temporal relevance, forming a high-dimensional interaction map. In the third phase, graph convolution is used to update node representations within each graph, allowing contextual information aggregation through learned adjacency matrices. Hyper-graph convolution extends this to include higher-order relationships across multiple graphs and temporal instances. In the fourth phase, a self-attention mechanism refines graph representations by dynamically adjusting the significance of nodes within and across modalities. This process focuses on the most relevant features for narrative generation. Finally, the event ratiocination generator uses the enriched feature representations to generate narratives using BART Lewis et al. (2019).
@@ 5.9 Discussion
Specifically, the first three techniques, presented in Sections 5.15.2, and 5.3 employed a CNN encoder for visual feature extraction and a RNN decoder for language generation. The methods detailed in Sections 5.1 and 5.3, offered a comparative analyses on the VIST dataset, with METEOR scores of 31.4 and 37.2, respectively. Notably, the latter technique surpasses the former, which can be attributed to its transitional training approach.
@@
Table 11 summarizes the performance of the methods described above, along with other methods from the literature that contributed to the field of VSG. The listed models range from earlier methods using CNNs and RNNs to more recent advances that integrate object detection models like Faster R-CNN Ren et al. (2015) and Mask-RCNN He et al. (2017), and LLMs like GPT-2 Radford et al. (2019) and BART Lewis et al. (2019). The first two methods by Park et al. (2015) Park and Kim (2015) and Park et al. (2019) Park et al. (2018) are not comparable to the other methods as they do not use the VIST dataset. The remaining methods are evaluated on the VIST dataset with the exception of CharGrid (2023) Hong et al. (2023) that is evaluated on the VWP dataset. The authors of CharGrid trained and evaluated the performance of TAPM on the VWP Hong et al. (2023) dataset, making the two methods directly comparable and it is likely that CharGrid would perform slightly better than TAPM on the VIST dataset. It is however not possible to know if CharGrid would perform better than HEGR on the VIST dataset without further experiments.
@@
Park et al. (2015) Park and Kim (2015) CNN, Bi-RNN RNN Blog Posts 3.49 - 8.78
Park et al. (2019) Park et al. (2018) Resnet-101/152, LSTM LSTM ActivityNet 9.91 - 16.48
Huang et al. (2015) Huang et al. (2016) CNN, GRU GRU VIST - - 31.4
Xu et al. (2015) Xu et al. (2015) VGGNet LSTM VIST - 28.94 32.98
Yu et al. (2017) Yu et al. (2017) ResNet-101, Bi-GRU GRU VIST - 29.53 34.12
CST (2018) Gonzalez-Rico and Pineda (2018) CNN, LSTM LSTM VIST 12.7 29.2 34.4
AREL (2018) Wang et al. (2018) ResNet-152, Bi-GRU GRU VIST 14.1 29.5 35.0
HSRL (2019) Huang et al. (2018) ResNet-152 LSTM VIST 12.32 30.84 35.23
GLACNet (2019) Kim et al. (2018) ResNet-152, Bi-LSTM LSTM VIST - - 30.63
StoryAnchor (2020) Modi and Parde (2019) ResNet-152, Bi-GRU GRU VIST 14.0 30.0 35.5
INET (2020) Jung et al. (2020) CNN, GRU GRU VIST 14.7 29.7 35.5
AOG-LSTM (2023) Liu et al. (2023) RestNet-152 LSTM VIST 12.9 30.1 36.0
CoVS (2023) Gu et al. (2023) RestNet-152, GRU GRU VIST 15.2 30.8 36.5
TAPM (2021) Yu et al. (2021) 1 Faster R-CNN, ResNet-152 GPT-2 s VIST - 33.1 37.2
HEGR (2021) Zheng et al. (2021) 1,2 Mask-RCNN, BERT BART VIST 16.4 37.4 37.8
TAPM (2021) Hong et al. (2023) Faster R-CNN, ResNet-152 GPT-2 VWP 5.19 25.09 32.38
CharGrid (2023) Hong et al. (2023)2 Mask-RCNN, SwinTransformer GPT-2 VWP 5.42 25.01 33.03
@@
Table 11: Analysis of Visual Story Generation works. The columns “B-4”, “R-L”, and “M” represent the BLEU-4, ROUGE-L, and METEOR scores, respectively. All works, except Huang et al. (2015) Huang et al. (2016) and CST (2018) Gonzalez-Rico and Pineda (2018), use attention mechanisms. All works use images as input, except ActivityNet Fabian Caba Heilbron and Niebles (2015) (it uses videos). The mark 1 indicates works that used extra training data (direct comparison may not be fair). The mark 2 indicates works that explicitly model some story elements, namely the characters by the detection of objects in the case of CharGrid (2023) and HEGR (2021) and the setting in the case of HEGR (2021).
6 Real-World Applications
As VSG technologies mature, they are increasingly being deployed in real-world scenarios. These applications demonstrate the practical potential of VSG systems and highlight new challenges that arise when deploying these technologies in dynamic, context-rich environments. This section explores emerging applications that integrate visual storytelling with contextual information, location awareness, and real-time data processing.
6.1 Location-Aware Visual Storytelling
Suwono et al. Suwono et al. (2023) introduced location-aware visual question generation with lightweight models, representing an important advancement in spatially-grounded narrative applications. Their work addresses the challenge of generating contextually relevant questions and narratives that are informed by geographical location, opening new possibilities for travel documentation, educational applications, and location-based social media platforms. The location-aware approach integrates geographical metadata with visual content analysis to create narratives that are visually grounded and spatially contextualized. This integration enables the generation of stories that can reference specific landmarks, cultural contexts, and geographical features that would be meaningful to users familiar with particular locations. The use of lightweight models addresses practical deployment constraints, making the technology suitable for mobile applications and resource-constrained environments.
6.2 Automotive and Real-Time Context Integration
Belz et al. Belz et al. (2024) explored story-driven approaches in automotive contexts, investigating how real-time contextual information can enhance automated storytelling systems. Their research represents a step toward practical deployment of VSG in everyday scenarios, particularly in automotive systems where visual storytelling could enhance the driving experience and provide meaningful documentation of journeys. The automotive application domain presents challenges and opportunities for visual storytelling. In-car camera systems continuously capture visual data, creating an opportunity for real-time narrative generation that could help drivers and passengers understand and remember their journeys. The integration of real-time context information—such as speed, location, weather conditions, and traffic patterns enables the generation of richer, more informative narratives that extend beyond visual description. Their work demonstrates how dynamic contextual data can be incorporated into storytelling algorithms, showing measurable improvements in narrative relevance and user engagement when real-time information is provided. This approach could lead to applications in autonomous vehicles, where passengers might receive automatically generated travel narratives, or in fleet management systems where journey documentation becomes an automated process.
7 Conclusions